OpenStudy (anonymous):

Having a really hard time remembering this... How many atoms of each element are in 5 moles of aluminum carbonate?

OpenStudy (aaronq):

Each formula unit of \(Al_2(CO_3)_3\) has 2 atoms of Al, 3 atoms of C and 9 of O so multiply the number of moles by the number of atoms in the formula unit, THEN multiply by avogadro's number, \(6.022∗10^{23}\)

OpenStudy (anonymous):

still having a bit of trouble with it. can you maybe explain it in a simpler way? @aaronq

OpenStudy (sasquat):

Id say use dimensional analysis to convert between the units you want. In this question you want to convert from atoms to moles. So first find the molar mass of aluminum carbonate and then multiply the number of moles of aluminum carbonate with the number of atoms: 6.022 x 10^23. Hope this helped :)

OpenStudy (aaronq):

In the "molecule" (called formula unit when talking about inorganic compounds), there are a number of atoms, these are indicated by the small numbers at the bottom right of each elemental symbol, \(\sf Al_\color{red}2(CO_\color{red}3)_\color{red}3\)- these numbers are called "subscripts", highlighted red here. in 1 formula unit of aluminium carbonate, there are 2 atoms of Al, 3 atoms of C, and 9 atoms of O now they're asking how many moles in 5 moles? 1 mole aluminium carbonate has 2 moles of Al, 3 moles of C and 9 moles of O. so 5 moles of aluminium carbonate have: 2*5= 10 moles of Al 3*5=15 moles of C 9*5=45 moles of O so to find atoms now, we use: \(\sf atoms=moles*avogadro's~number\) \(\sf Al ~atoms=10 ~moles*6.022*10^{23}~atoms/mole=3.011*10^{24}~atoms\) so the same for the rest

OpenStudy (anonymous):

you can watch a tylerdewitt vedio on this topic